OpenStudy (anonymous):

What is the function form of 5x-y=10?

OpenStudy (jdoe0001):

well, function form? http://www.mathsisfun.com/algebra/standard-form.html ?

OpenStudy (anonymous):

5x-y=10 -y=10-5x y=5x-10 It's a line with inclination 5 and y intercept point -10

OpenStudy (anonymous):

and x intercept in 2

OpenStudy (anonymous):

Thanks